OpenStudy (eleanorbrit11488):

There are 60 students in a speech contest. Yesterday, 2/3 of them gave their speeches. Today, 1/5 of the remaining students gave their speeches. How many students still haven't given their speeches?

Vocaloid (vocaloid):

hint: what is 2/3 of 60?

OpenStudy (eleanorbrit11488):

40?

OpenStudy (mathmale):

Right. Evaluate (2/3)*60. Then for today, mult. your result by 1/5.

OpenStudy (eleanorbrit11488):

8?

Vocaloid (vocaloid):

good, so the number left is 60 - 40 = 20 then multiply 20 by 1/5

OpenStudy (mathmale):

Yes. So, how many students, total, have already given their speeches? How many have not?

OpenStudy (eleanorbrit11488):

52 have and 8 have not?

OpenStudy (mathmale):

I'm afraid we're going to have to backtrack. Yesterday 2/3 rds of the 60 students gave their speeches, leaving how many who had not? It's not 40. Think thru this, please.

OpenStudy (mathmale):

Please refer to Vocaloid's most recent post. It's correct.

OpenStudy (mathmale):

@skull patrol: Your response and vocaloid's disagree. Please work this out directly with Vocaloid.

OpenStudy (mathmale):

The issue here is: After yesterday, how many students had NOT yet given their speeches?

OpenStudy (mathmale):

"Vocaloid Mathlete Best Response Medals 1 good, so the number left is 60 - 40 = 20 then multiply 20 by 1/5"

OpenStudy (skullpatrol):

There are 60 students in a speech contest. Yesterday, 2/3 of them gave their speeches. $$2/3(60) = 40$$ Today, 1/5 of the remaining students gave their speeches. $$60-40 = 20$$ $$1/5(20) = 4$$ How many students still haven't given their speeches? $$60 – 40 – 4 = ?$$

OpenStudy (mathmale):

Nice work. But I have also recognized Vocaloid for having suggested this approach earlier. @eleanorbrit11488 : Can you now finish this problem?